Calculate the pH of a solution which contains 9.9 ml of 1 M HCI and 100 ml of 0.1 M NaOH

Grade:10

1 Answers

Arun
6 years ago
Na OH reacts with H Cl : 
Na OH + H Cl → Na Cl  + H2O 
1 mol Na OH reacts with 1 mol HCl 

Mol H Cl in 9 . 9 m L of 1 . 0 M solution =  9.9 /1000 * 1.0  =  0.0099 mol H Cl 
Mol Na OH in 100mL of  0.1 M solution = 100 / 1000 * 0.1  =  0.01 mol Na OH 
On mixing the 0.0099 mol H Cl will be neutralized and 0.01 -  0.0099  =  1 * 10 ^ - 4  mol Na OH dissolved in 109.9 m L solution 
Morality of Na OH solution  =  ( 1 * 10 ^ - 4 )  /  0.1099  =  9.099 * 10 ^ - 4 M 
p OH = - log ( 9.099 * 10 ^ -  4) 
p OH  = 3.04 
pH  =  14.00  – p OH 
pH  = 14.00  -  3.04 
pH  =  10.96
 
Regards
Arun